OpenStudy (anonymous):

WILL FAN AND MEDAL! Write each equation in standard form. y-1/2=x+1/2

OpenStudy (anonymous):

Standard form of equation = \[Ax + By = C\]

OpenStudy (anonymous):

Why does the author of "Enrico Fermi: The 'Italian Navigator'" compare Dr. Fermi to a ship's captain? A. because Fermi carefully worked through unknown and potentially dangerous territory B. because Fermi worked with a crew of people who would not cooperate with him C. because Fermi's experiments occurred in the captain's quarters on a ship D. because Fermi was "lost at sea" in his project

OpenStudy (anonymous):

Dude this is my question

OpenStudy (anonymous):

my bad?

OpenStudy (anonymous):

whats up with that

OpenStudy (anonymous):

@Lily7 please help me ill medal you

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

WHY AM I SO STUPID TORWARDS THIS

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

@wio please help

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You are not stupid. y-1/2=x+1/2 Ax +By = C y= x+1/2 +1/2 -x + y = 2/2 => -x +y = 1

OpenStudy (anonymous):

r u sure thats right?

OpenStudy (anonymous):

I am

OpenStudy (anonymous):

ok